OpenStudy (anonymous):

Evaluate the function x=2: f(x)= 7x-5. A,9. B,19 c,-9 d,-19

OpenStudy (trojanpoem):

Plug x in the function f(2) = 14- 5 = 9

OpenStudy (anonymous):

What's that mean ? @trojanpoem

OpenStudy (trojanpoem):

Evaluate the function at x = 2 you set each x to 2 in y = f(x) = 7x - 5 to find y

OpenStudy (anonymous):

So the answer is 9??! If I'm correct @trojanpoem

OpenStudy (trojanpoem):

Correct.

OpenStudy (anonymous):

Thanks a lot !!! I'm new to openstudy so Idk how to medal u😂but can I keep asking you questions ?

OpenStudy (anonymous):

@trojanpoem

OpenStudy (trojanpoem):

sure.

OpenStudy (anonymous):

Candy is on sale for $0.75 each. You have a coupon for $0.25 off your total purchase.write a funtionc rule for the cost of n pieces of candy

OpenStudy (anonymous):

A,c(n) = 0.25n -0.75 @trojanpoem

OpenStudy (anonymous):

B,c(n)= 0.25n -0.75

OpenStudy (trojanpoem):

c(n) = 0.75n - 0.25

OpenStudy (anonymous):

C,c(n)=1.0n

OpenStudy (trojanpoem):

what's d ?